OpenStudy (anonymous):

The probability that a man will hit a target is 2 / 3. If he shoots at the target until he hits it for the first time, find the probability that it will take him 5 shots to hit the target

OpenStudy (anonymous):

@Directrix @jim_thompson5910

OpenStudy (anonymous):

this one is a hard one

jimthompson5910 (jim_thompson5910):

This is basically asking the question of "what is the probability that he will hit the target on the 6th try" This is assuming that once he hits the target, he can no longer shoot it

jimthompson5910 (jim_thompson5910):

So you would use the geometric distribution to answer this question

jimthompson5910 (jim_thompson5910):

which means you use this formula P(X = k) = (1-p)^(k-1)*p where p = probability of success of each individual and independent trial k = trial number where the first success is encountered

OpenStudy (anonymous):

i know the book says its 2 / 243, but i dont know process

jimthompson5910 (jim_thompson5910):

you just plug in p = 2/3 and k = 6 evaluate and simplify

OpenStudy (anonymous):

got it. thmks

jimthompson5910 (jim_thompson5910):

oh wait, k = 5, not 6 but everything else is still the same